A certain pendulum has a period of 1.7s on Earth. What is the period of this pendulum on a planet that has a gravitational acceleration of 15m/s 2. 


1
Expert's answer
2022-02-10T09:18:13-0500

Given:

"T_E=1.7\:\rm s"

"g_E=9.8\rm \: m/s^2"

"g_P=15\rm \: m/s^2"


The period of oscillation of a simple pendulum

"T=2\pi\sqrt{\frac{l}{g}}"

Hence

"T_P=T_E\sqrt{\frac{g_E}{g_P}}=1.7\sqrt{\frac{9.8}{15}}=1.4\:\rm s"